OpenStudy (anonymous):

is this right : y^2-4x=4 : x=-1+y^2/4

OpenStudy (curry):

i got this one

OpenStudy (curry):

ye tis right

OpenStudy (anonymous):

\[y^2 - 4x = 4\]\[\implies y^2 = 4x + 4\]\[\implies y^2 - 4 = 4x\]\[\implies x = \frac{y^2}{4} - 1\]

OpenStudy (curry):

polpak can you help me on my question